OpenStudy (anonymous):

Im horrible at word problems. Calc help plz :/ My aunt and I were approaching the same intersection, she from the south and I from the west. She was traveling at a steady speed of 5 miles/hour, while I was approaching the intersection at 50 miles/hour. At a certain instant in time, I was three-tenths of a mile from the intersection, while she was one-twentieth of a mile from it. How fast were we approaching each other at that instant? (Give your answer to the nearest integer.) answer should be in MPH

OpenStudy (anonymous):



OpenStudy (anonymous):

we are going from the west , so we have distance from intersection as i labeled "x" and you are told that \[x'=50\] (well actually it should be -50 because x is getting smaller) likewise grandma's distance from intersection is y and you are told \[y'=5\]

OpenStudy (anonymous):

if we call the distance we are from each other "z" then pythagoras tells us that \[x^2+y^2=z^2\] differentiating gives us \[2xx'+2yy'=2zz'\] all the numbers are known except \[z'\] which is the one we want, so solve for that \[x'=50,y'=5, x =.3, y = .05, z = \sqrt{.3^2+.05^2}\] plug them in and solve for \[z'\]